# Physics meets Topology: Physics-informed topological neural networks for learning rigid body dynamics

### v1

@@ 5.2 Physics-informed message passing
B. Processing inter-object collisions using mesh faces. This second major part, particularly critical in our framework as it models the inter-object collisions, is presented in Fig. 9b and 9c. To ensure energy conservation during each individual collision between two mesh triangles $x_{s}^{(2)}$ and $x_{r}^{(2)}$, the messages $m^{2\rightarrow 2}_{s\rightarrow r}$ and $m^{2\rightarrow 2}_{r\rightarrow s}$ are built in a permutation-invariant manner using the unidirectional collision contacts cells $x^{(3)}_{s\rightarrow r}$ and $x^{(3)}_{r\rightarrow s}$. First, the collision effect from $x^{(2)}_{s}$ on $x^{(2)}_{r}$ is computed using $x^{(3)}_{s\rightarrow r}$ (step 2, Eq. 10). Then, the new embedding $\bm{h}^{\prime\prime(2)}_{r}$ of the receiver triangle $x^{(2)}_{r}$ is calculated by aggregating all incoming collision messages on this mesh triangle (step 3, Eq. 11), leading to:
$\bm{m}_{s}^{2\rightarrow 3}=\text{MLP}^{\text{processor}}_{m2\rightarrow 3}(\bm{h}_{s}^{\prime(2)})\qquad\bm{h}^{\prime(3)}_{s\rightarrow r}=\text{MLP}^{\text{processor}}_{3}([\bm{h}_{s\rightarrow r}^{(3)},\bm{m}_{s}^{2\rightarrow 3},\bm{m}_{r}^{2\rightarrow 3}])$ (10)
@@
$\bm{m}_{s\rightarrow r}^{3\rightarrow 2}=\text{MLP}^{\text{processor}}_{m3\rightarrow 2}(\bm{h}_{s\rightarrow r}^{\prime(3)})\qquad\bm{h}^{\prime\prime(2)}_{r}=\text{MLP}^{\text{processor}}_{2^{\prime}}([\bm{h}_{r}^{\prime(2)},\sum_{\forall j\in\mathcal{N}_{\nearrow,1}(x^{(2)}_{r})}\bm{m}_{j\rightarrow r}^{3\rightarrow 2}])$ (11)
@@
where $\bm{m}_{s}^{2\rightarrow 3}$ and $\bm{m}_{r}^{2\rightarrow 3}$ are learnable messages from sender triangle $s$ and receiver triangle $r$, $\bm{h}^{\prime(3)}_{s\rightarrow r}$ is the updated collision contact embedding after message-passing step 2, $\bm{m}_{s\rightarrow r}^{2\rightarrow 3}$ is a learnable message from the collision contact $x^{(3)}_{s\rightarrow r}$ to the receiver triangle $x^{(2)}_{r}$, and $\bm{h}^{\prime\prime(2)}$ are the new mesh triangle embeddings after message-passing step 3.
@@
C. Updating intra-objects cells after collisions. This part handles the intra-object updates by integrating the effects of incoming collisions and the current velocity and momentum. The information about incoming collisions is embedded in the updated triangle features $\bm{h}^{\prime\prime(2)}$, while the velocity and momentum at time $t$ are encoded in the node embeddings $\bm{h}^{(0)}$ and object embeddings $\bm{h}^{(4)}$. In step 4 (Eq. 12), updated object embeddings $\bm{h}^{\prime(4)}$ are computed using the incoming collisions. Subsequently, the final object embeddings $\bm{h}^{\prime\prime(4)}$ and node embeddings $\bm{h}^{\prime(0)}$ are derived (Eq. 13-15), where $k$ is the index of the object containing node $i$ and $N$ is the total number of nodes in object $k$.
@@
$\bm{m}_{i}^{2\rightarrow 4}=\text{MLP}^{\text{processor}}_{m2\rightarrow 4}(\bm{h}_{i}^{\prime\prime(2)})\qquad\bm{h}^{\prime(4)}_{i}=\text{MLP}^{\text{processor}}_{4}([\bm{h}_{i}^{(4)},\sum_{\forall j\in\mathcal{N}_{\searrow,2}(x^{(4)}_{i})}\bm{m}_{j}^{2\rightarrow 4}])$ (12)
$\bm{m}_{i}^{0\rightarrow 4}=\text{MLP}^{\text{processor}}_{m0\rightarrow 4}(\bm{h}_{i}^{(0)})\qquad\bm{m}_{i}^{4\rightarrow 0}=\text{MLP}^{\text{processor}}_{m4\rightarrow 0}(\bm{h}_{i}^{\prime\prime(4)})$ (13)
$\bm{h}^{\prime(0)}_{i}=\text{MLP}^{\text{processor}}_{0}([\bm{h}_{i}^{(0)},\bm{m}_{k}^{4\rightarrow 0}])$ (14)
@@
$\bm{h}^{\prime\prime(4)}_{n}=\text{MLP}^{\text{processor}}_{4^{\prime}}([\bm{h}_{n}^{(0)},\frac{1}{N}\sum_{\forall j\in\mathcal{N}_{\searrow,4}(x^{(4)}_{n})}\bm{m}_{j}^{0\rightarrow 4}])$ (15)
@@
where $\bm{m}^{2\rightarrow 4}$ are learnable messages from mesh triangles to their parent object, $\bm{h}_{i}^{\prime(4)}$ are the new object embeddings after message-passing step 4, $\bm{m}^{0\rightarrow 4}$ are learnable messages from nodes to object cells, $\bm{m}^{4\rightarrow 0}$ are learnable messages from object to node cells, $\bm{h}^{\prime(0)}$ the final node embeddings, and $\bm{h}^{\prime\prime(4)}$ the final object embeddings after message passing step 5.
@@
D. Predicting accelerations and the next pose. This last part contains the final decoding step of our Encode-Process-Decode architecture (step 6). It simply consists of a readout step on the object embeddings $\bm{h}^{\prime\prime(4)}$ and node embeddings $\bm{h}^{\prime(0)}$ with two MLPs (Eq. 16).
$\hat{\ddot{\bm{x}}}_{i}=\text{MLP}^{\text{decoder}}_{0}(\bm{h}_{i}^{\prime(0)})\qquad\hat{\ddot{\bm{x}}}_{n}=\text{MLP}^{\text{decoder}}_{4}(\bm{h}_{i}^{\prime\prime(4)})$ (16)
@@
where $\hat{\ddot{\bm{x}}}_{i}$ and $\hat{\ddot{\bm{x}}}_{n}$ are the predicted accelerations for nodes and objects, respectively.
@@
Finally, the last element of our framework consists in predicting the poses of all objects at time $t+1$. It is exclusively applied during inference, not training. Here, we use the second-order forward-Euler integration to predict the next node positions $\hat{\bm{x}}^{t+1}_{i}$. To ensure the objects maintain their rigidity over extended rollouts, we use shape_matching . Using the predicted node positions $\{\hat{\bm{x}}_{i}\}_{i=1..N}$ of each object $k$, we compute the updated center of mass $\hat{\bm{x}}^{t+1}_{k}$ and quaternion orientation $\hat{\bm{q}}^{t+1}_{k}$ for each object. Finally, we apply the rotation $\hat{\bm{q}}^{t+1}_{k}$ then translation $\hat{\bm{x}}^{t+1}_{k}$ to the reference mesh $M$, and obtain the final node positions predictions.
@@ 5.3 Performance evaluation metrics
We use the Root Mean Square Error (RMSE) between the predicted and ground truth object trajectories, on both position and rotation. We report the average RMSE per object in meters (for position) and degrees (for rotation). The exact formulas are given in Eq. 17 and 18, where $\langle\bm{q}\rangle$ is the norm of the vector part of quaternion $\bm{q}$ and $\times$ is the Hamilton product.
$\text{RMSE}^{\text{pos}}=\sqrt{\frac{1}{K}\sum_{k=1}^{K}\lVert\hat{\bm{x}}-\bm{x}\rVert^{2}}$ (17)
$\text{RMSE}^{\text{ori}}=\sqrt{\frac{1}{K}\sum_{k=1}^{K}\frac{360}{\pi}\arcsin{(\langle\hat{\bm{q}}\times\bm{q}^{-1}\rangle)}}$ (18)
@@
The datasets used for the experiments have been generated with the Kubric library kubric . The exact parameters used to generate the datasets are given in Supplementary Information Sec. 2.
@@
The implementation of the proposed method is based on PyTorch pytorch and TopoNetX toponetx . The source code and model checkpoints will be made publicly available on GitHub after acceptance.
@@ Acknowledgments
We thank Vinay Sharma and Ismail Nejjar for helpful discussions. This study was supported by the Swiss National Science Foundation (grant number 200021_200461).
@@
Additional information
Correspondence and requests for materials should be addressed to Olga Fink.

### v3

@@ 5.2.2 Processing inter-object collisions using mesh faces
This second major part, particularly critical in our framework as it models the inter-object collisions, is presented in Figure 9b and 9c. To ensure energy conservation during each individual collision between two mesh triangles $x_{s}^{(2)}$ and $x_{r}^{(2)}$, the messages $m^{2\rightarrow 2}_{s\rightarrow r}$ and $m^{2\rightarrow 2}_{r\rightarrow s}$ are built in a permutation-invariant manner using the unidirectional collision contacts cells $x^{(3)}_{s\rightarrow r}$ and $x^{(3)}_{r\rightarrow s}$. First, the collision effect from $x^{(2)}_{s}$ on $x^{(2)}_{r}$ is computed using $x^{(3)}_{s\rightarrow r}$ (step 2, Eq. 10). Then, the new embedding $\mathbf{h}^{\prime\prime(2)}_{r}$ of the receiver triangle $x^{(2)}_{r}$ is calculated by aggregating all incoming collision messages on this mesh triangle (step 3, Eq. 11), leading to:
$\mathbf{m}_{s}^{2\rightarrow 3}=\text{MLP}^{\text{processor}}_{m2\rightarrow 3}(\mathbf{h}_{s}^{\prime(2)})\qquad\mathbf{h}^{\prime(3)}_{s\rightarrow r}=\text{MLP}^{\text{processor}}_{3}([\mathbf{h}_{s\rightarrow r}^{(3)},\mathbf{m}_{s}^{2\rightarrow 3},\mathbf{m}_{r}^{2\rightarrow 3}])$ (10)
@@
$\mathbf{m}_{s\rightarrow r}^{3\rightarrow 2}=I(\mathbf{h}_{s\rightarrow r}^{\prime(3)})\qquad\mathbf{h}^{\prime\prime(2)}_{r}=\text{MLP}^{\text{processor}}_{2^{\prime}}([\mathbf{h}_{r}^{\prime(2)},\sum_{\forall j\in\mathcal{N}_{\nearrow,1}(x^{(2)}_{r})}\mathbf{m}_{j\rightarrow r}^{3\rightarrow 2}])$ (11)
@@
where $I$ is the identity function, $\mathbf{m}_{s}^{2\rightarrow 3}$ and $\mathbf{m}_{r}^{2\rightarrow 3}$ are learnable messages from sender triangle $s$ and receiver triangle $r$, $\mathbf{h}^{\prime(3)}_{s\rightarrow r}$ is the updated collision contact embedding after message-passing step 2, $\mathbf{m}_{s\rightarrow r}^{2\rightarrow 3}$ is a learnable message from the collision contact $x^{(3)}_{s\rightarrow r}$ to the receiver triangle $x^{(2)}_{r}$, and $\mathbf{h}^{\prime\prime(2)}$ are the new mesh triangle embeddings after message-passing step 3.
5.2.3 Updating intra-objects cells after collisions
@@
This part handles the intra-object updates by integrating the effects of incoming collisions and the current velocity and momentum. The information about incoming collisions is embedded in the updated triangle features $\mathbf{h}^{\prime\prime(2)}$, while the velocity and momentum at time $t$ are encoded in the node embeddings $\mathbf{h}^{(0)}$ and object embeddings $\mathbf{h}^{(4)}$. In step 4 (Eq. 12), updated object embeddings $\mathbf{h}^{\prime(4)}$ are computed using the incoming collisions. Subsequently, the final object embeddings $\mathbf{h}^{\prime\prime(4)}$ and node embeddings $\mathbf{h}^{\prime(0)}$ are derived (Eq. 13-15), where $k$ is the index of the object containing node $i$, $F$ is the total number of faces in object $k$, and $N$ is the total number of nodes in object $k$.
@@
$\mathbf{m}_{i}^{2\rightarrow 4}=\text{MLP}^{\text{processor}}_{m2\rightarrow 4}(\mathbf{h}_{i}^{\prime\prime(2)})\qquad\mathbf{h}^{\prime(4)}_{k}=\text{MLP}^{\text{processor}}_{4}([\mathbf{h}_{k}^{(4)},\frac{1}{F}\sum_{\forall j\in\mathcal{N}_{\searrow,2}(x^{(4)}_{k})}\mathbf{m}_{j}^{2\rightarrow 4}])$ (12)
$\mathbf{m}_{i}^{0\rightarrow 4}=\text{MLP}^{\text{processor}}_{m0\rightarrow 4}(\mathbf{h}_{i}^{(0)})\qquad\mathbf{m}_{k}^{4\rightarrow 0}=\text{MLP}^{\text{processor}}_{m4\rightarrow 0}(\mathbf{h}_{k}^{\prime(4)})$ (13)
$\mathbf{h}^{\prime(0)}_{i}=\text{MLP}^{\text{processor}}_{0}([\mathbf{h}_{i}^{(0)},\mathbf{m}_{k}^{4\rightarrow 0}])$ (14)
@@
$\mathbf{h}^{\prime\prime(4)}_{k}=\text{MLP}^{\text{processor}}_{4^{\prime}}([\mathbf{h}_{k}^{\prime(4)},\frac{1}{N}\sum_{\forall j\in\mathcal{N}_{\searrow,4}(x^{(4)}_{k})}\mathbf{m}_{j}^{0\rightarrow 4}])$ (15)
@@
where $\mathbf{m}^{2\rightarrow 4}$ are learnable messages from mesh triangles to their parent object, $\mathbf{h}^{\prime(4)}$ are the new object embeddings after message-passing step 4, $\mathbf{m}^{0\rightarrow 4}$ are learnable messages from nodes to object cells, $\mathbf{m}^{4\rightarrow 0}$ are learnable messages from object to node cells, $\mathbf{h}^{\prime(0)}$ the final node embeddings, and $\mathbf{h}^{\prime\prime(4)}$ the final object embeddings after message passing step 5.
5.2.4 Predicting accelerations and the next pose
@@
This last part contains the final decoding step of our Encode-Process-Decode architecture (step 6). It simply consists of a readout step on the object embeddings $\mathbf{h}^{\prime\prime(4)}$ and node embeddings $\mathbf{h}^{\prime(0)}$ with two MLPs (Eq. 16).
$\hat{\ddot{\mathbf{x}}}_{i}=\text{MLP}^{\text{decoder}}_{0}(\mathbf{h}_{i}^{\prime(0)})\qquad\hat{\ddot{\mathbf{x}}}_{k}=\text{MLP}^{\text{decoder}}_{4}(\mathbf{h}_{k}^{\prime\prime(4)})$ (16)
@@
where $\hat{\ddot{\mathbf{x}}}_{i}$ and $\hat{\ddot{\mathbf{x}}}_{k}$ are the predicted accelerations for nodes and objects, respectively.
@@
Finally, the last element of our framework consists in predicting the poses of all objects at time $t+1$. It is exclusively applied during inference, not training. Here, we use the second-order forward-Euler integration to predict the next node positions $\hat{\mathbf{x}}^{t+1}_{i}$. To ensure the objects maintain their rigidity over extended rollouts, we use shape matching shape_matching . Using the predicted node positions $\{\hat{\mathbf{x}}_{i}\}_{i=1..N}$ of each object $k$, we compute the updated center of mass $\hat{\mathbf{x}}^{t+1}_{k}$ and quaternionic orientation $\hat{\mathbf{q}}^{t+1}_{k}$ for each object. Finally, we apply the rotation $\hat{\mathbf{q}}^{t+1}_{k}$ then translation $\hat{\mathbf{x}}^{t+1}_{k}$ to the reference mesh $M^{t}$, and obtain the final node positions predictions.
The choice of aggregation function in HOPNet’s message passing depends on whether the number of contributing entities is fixed or variable. When the count is constant—as in the case of updating face features using information from nodes, edges, and object (Eq. 9)—we use SUM, enabling the network to learn appropriate scaling factors. In contrast, when the number of messages varies, such as during information exchange between nodes and objects (Eq. 12 and 15), we use MEAN to maintain consistency across different mesh resolutions. An exception is made during collision processing (Eq. 11), where we apply SUM despite the varying number of incoming collisions per face, in order to preserve Newton’s law of force summation.
HOPNet’s physics-informed message-passing strategy is partially illustrated in Figure 10, which depicts parts A and B of the message-passing. A comprehensive overview of the full message-passing strategy is provided in Supplementary Figure 3 (Suppl. Sec. 2.2).
@@ 5.3 Performance evaluation metrics
We use the Root Mean Square Error (RMSE) between the predicted and ground truth object trajectories, on both position and rotation. We report the average RMSE per object in meters (for position) and degrees (for rotation). The exact formulas are given in Equations 17 and 18, where $\langle\mathbf{q}\rangle$ is the norm of the vector part of quaternion $\mathbf{q}$ and $\times$ is the Hamilton product.
$\text{RMSE}^{\text{pos}}=\sqrt{\frac{1}{K}\sum_{k=1}^{K}\lVert\hat{\mathbf{x}}-\mathbf{x}\rVert^{2}}\hskip-14.22636pt$ (17)
$\text{RMSE}^{\text{ori}}=\sqrt{\frac{1}{K}\sum_{k=1}^{K}\frac{360}{\pi}\arcsin{(\langle\hat{\mathbf{q}}\times\mathbf{q}^{-1}\rangle)}}\hskip 0.0pt$ (18)
5.4 Ethics and inclusion statement
This work introduces a general-purpose, learning-based framework for simulating rigid body dynamics across diverse environments, with applications in fields such as robotics, mechanics, and material science. By learning directly from data, our approach can model physical interactions in complex or partially observed environments, supporting more inclusive and adaptable simulation tools.
The models are trained on synthetic datasets that do not involve human subjects or sensitive attributes. We encourage responsible deployment of this technology, especially in scenarios involving real-world decision-making or safety-critical applications.
@@ 5.4 Ethics and inclusion statement
The data generated in this study for training and testing have been deposited in the Zenodo database under accession code https://doi.org/10.5281/zenodo.15800434. The datasets used in our experiments were generated using the open-source Kubric library kubric . The full datasets can be exactly reproduced using the configuration provided in Supplementary Information Section 2.
@@
The implementation of the proposed method is based on PyTorch pytorch and TopoNetX toponetx . Our reproducible code is available at https://github.com/AmauryWEI/hopnet.
@@ Acknowledgments
We thank Vinay Sharma and Ismail Nejjar for helpful discussions. This work was supported by the Swiss National Science Foundation (SNSF) Grant Number 200021_200461.